# Charged spherically symmetric and slowly rotating charged black hole solutions in bumblebee gravity

### v1

@@ Abstract
In this paper, we present charged spherically symmetric black hole solutions and slowly rotating charged solutions in bumblebee gravity with and without a cosmological constant. The static spherically symmetric solutions describe the Reissner-Nordström-like black hole and Reissner-Nordström-(anti) de Sitter-like black hole, while the stationary and axially symmetric soltuions describe Kerr-Newman-like black hole and Kerr-Newman-(anti) de Sitter-like black hole. We utilize the Hamilton-Jacobi formalism to study the shadows of the black holes. Additionally, we investigate the effect of the electric charge and Lorentz-violating parameters on the radius of the shadow reference circle and the distortion parameter. We find that the radius of the reference circle decreases with the Lorentz-violating parameter and charge parameter, while the distortion parameter increases with the Lorentz-violating parameter and the charge parameter.
@@ I INTRODUCTION
Several quantum gravity theories have been proposed Birrell and Davies (1984); Maldacena (1998); Aharony et al. (2000); Gubser et al. (1998); Alfaro et al. (2000, 2002); Rovelli and Smolin (1990), but it is needed the experiments at the Planck scale ($\sim 10^{19}$ GeV) to observe the quantum effect, this is beyond the ability of current experiments. However, some quantum gravity theories assume that Lorentz symmetry might be broken in the gravitational UV regime which is at our presently accessible low-energy scales. Stemming from the concept of spontaneous Lorentz symmetry breaking (LSB) in string theory Kostelecky and Samuel (1989a); Kostelecky (2004), the Standard Model extension was introduced as a suitable framework for violating Lorentz symmetry Colladay and Kostelecky (1997, 1998); Kostelecky (2004). The Standard Model extension has been widely studied Colladay and Kostelecky (2001); Kostelecky and Lehnert (2001); Kostelecky and Mewes (2001); Colladay and McDonald (2009); Berger et al. (2016); Carroll et al. (1990); Andrianov and Soldati (1995, 1998); Lehnert and Potting (2004); Baeta Scarpelli (2012); Brito et al. (2013).
@@
In 1989, Kostelecky and Samuel introduced the prototype of the bumblebee model Kostelecky and Samuel (1989b, a), a string-inspired framework featuring tensor-induced spontaneous Lorentz symmetry breaking. In the framework of bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B^{\mu}B_{\mu})$ acting on an vector field $B_{\mu}$ Kostelecky (2004); Kostelecky and Tasson (2011). Not just in Minkowski’s spacetime, The bumblebee model can also be explored in various spacetime geometries, including Riemann, Riemann-Cartan spacetimes Bailey and Kostelecky (2006); Maluf et al. (2013); Li and Övgün (2020); Bluhm et al. (2008a); Escobar and Martín-Ruiz (2017); Maluf et al. (2015); Hernaski (2014); Carroll et al. (2009).
@@
The pursuit of black hole solutions holds significance for any theory of gravity. In 2017, Casana et al. obtained an exact Schwarzschild-like black hole solution in the bumblebee gravity Casana et al. (2018). Subsequently, researchers have independently proposed exact solutions for traversable wormholes Övgün et al. (2019), Schwarzschild-anti-de Sitter (AdS)-like black holes Maluf and Neves (2021), slowly rotating Kerr-like black holes Ding et al. (2020), and other solutions Santos et al. (2015); Jha and Rahaman (2021); Filho et al. (2023); Xu et al. (2023); Ding et al. (2023) within bumblebee gravity. Moreover, in other scenarios within the framework of the Standard Model Extension, such as when considering the nonminimally coupled Kalb-Ramond field with a nonzero vacuum expectation value, certain black hole solutions have also been obtained Lessa et al. (2020); Kumar et al. (2020); Yang et al. (2023); Duan et al. (2023); Do and Kao (2020); Liu et al. (2024a). And the properties of the blakc holes were also studied Guo et al. (2023); Du et al. (2024); Kuang and Övgün (2022); Oliveira et al. (2019); Liu et al. (2023); Gomes et al. (2020); Kanzi and Sakallı (2019); Güllü and Övgün (2022); Oliveira et al. (2021); Kanzi and Sakallı (2021); Hosseinifar et al. (2024); Liu et al. (2024b, a).
@@ II EINSTEIN-BUMBLEBEE THEORY
$T^{M}_{\mu\nu}=\frac{1}{2\kappa}\left[(1+\gamma b^{2})(2F_{\mu\alpha}F_{\nu}^{\alpha}-\frac{1}{2}g_{\mu\nu}F^{\alpha\beta}F_{\alpha\beta})+\gamma B_{\mu}B_{\nu}F^{\alpha\beta}F_{\alpha\beta}\right].$ (7)
@@ III.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
where $\lambda$ is a constant.
@@
$\frac{d}{dr}[r^{2}\phi^{\prime}(r)]$ $=$ $0,$ (28)
by simplifying Eq. (24) and Eq. (25), we obtain:
@@
Here $\Sigma$ represents a 3-dimensional spacelike region with the induced metric $\gamma_{ij}^{(3)}$, the boundary $\partial\Sigma$ is a two-sphere located at spatial infinity with the induced metric $\gamma_{ij}^{(2)}=r^{2}{\leavevmode\nobreak\ d}\Omega^{2}$, while $n_{\mu}=(1,0,0,0)$ and $\sigma_{\mu}=(0,1,0,0)$ denote the unit normal vectors associated with $\partial\Sigma$ and $\Sigma$, respectively. The solution of $A(r)$ can also be rewritten as
@@
This solution is similar to the Reissner-Nordström (RN) solution. As $\ell$ tends to 0, it recovers to the RN solution. Additionally, it is evident that the metric functions approach to $A(r)\rightarrow 1$ and $S(r)\rightarrow 1+\ell$ at infinity, indicating that the spacetime is not asymptotically Minkowski. Similar to the Reissner-Nordström black hole, this solution exhibits two horizons:
@@
From this we can see that, the singularity only appears at $r=0$ (assuming $0<\left|\ell\right|\ll 1$). This is similar to the RN black hole.
@@ III.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
where $\lambda$ is a Lagrange multiplier Bluhm et al. (2008b). Note that the equation of motion for the Lagrange multiplier ensures the vacuum condition $X=0$, which also ensures that $V=0$ for any field $\lambda$ on-shell. However, unlike the vacuum condition in Case A, $V^{\prime}\neq 0$ is valid when the field $\lambda$ is nonzero. Thus, the potential $V$ contributes to the field equations. The constraint for the Lagrange multiplier $\lambda$ is providedthe field Eqs (10) and (11) .
@@
An analytical solution is be possible if and only if
@@
$A(r)$ $=$ $1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)r^{2}\Lambda_{e},$ (46)
$S(r)$ $=$ $\frac{1+\ell}{1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)r^{2}\Lambda_{e}},$ (47)
@@
As $\ell$ approaches to 0, this solution becomes RN-(A)ds black hole solution. Similarly, we compute the Kretschmann scalar in order to examine the singularity study the singularity. The expression for $K$ is
@@ IV.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
Ding et al. obtained a slowly rotating black hole solution within bumblebee gravity  Ding et al. (2020); Ding and Chen (2021). Following their approach, we obtained a slowly rotating charged black hole solution within bumblebee gravity. The metric in Boyer-Lindquist coordinates reads:
@@
When $\ell\rightarrow 0$, it recovers to the usual Kerr-Newman metric and when $a\rightarrow 0$, it becomes to metric (13) with Eqs. (30) and (31). By solving $\Delta_{r}=0$, we can obatin the radius of the black hole’s event horizon
$r_{\pm}=M\pm\sqrt{M^{2}-\frac{2(\ell+1)Q_{0}^{2}+(\ell+2)a^{2}}{\ell+2}}.$ (58)
It becomes evident that the mass and spin parameters of the black hole must satisfy.
$\frac{2(\ell+1)Q_{0}^{2}+(\ell+2)a^{2}}{\ell+2}\leq M^{2}.$ (59)
@@
$\Delta_{\mu\nu}=$ $R_{\mu\nu}$ $-\kappa\left[V^{\prime}\left(2B_{\mu}B_{\nu}+b^{2}g_{\mu\nu}\right)+B_{\mu}^{\ \alpha}B_{\nu\alpha}+Vg_{\mu\nu}\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}\right]$ (60)
@@
$\Delta_{\mu\nu}=\left\{\begin{aligned} \mathcal{O}(a^{2}\ell)&,&\quad\mu\nu=tt,rr,\phi\phi,\theta\theta,r\theta,t\phi\\ 0&,&\quad\text{otherwise},\end{aligned}\right.$ (61)
and the remaining components of $\bar{R}_{\mu\nu}=0$ are all zero. At the same time, we derive the equations of motion for the bumblebee field and the electromagnetic field. Only the $r$ component nonvanishing:
@@
Combining the above equations, we find that the rotating solution needs to satisfy the condition $\mathcal{O}(a^{2}\ell)$ being a small quantity. This implies that when considering this solution,$a$ and $\ell$ cannot both take large values simultaneously. Thus, for our solution, we need the condition that $a^{2}\ell$ is small quantities ($\frac{a^{2}\ell}{M^{2}}\ll 1$) for finite $Q_{0}$ , respectively.

### v2

@@ Abstract
In this paper, we study the scenario in which the matter field is an electromagnetic field nonminimally coupled to the bumblebee vector field. We present exact charged spherically symmetric black hole solutions and slowly rotating charged solutions in bumblebee gravity with and without a cosmological constant. The static spherically symmetric solutions describe the Reissner-Nordström-like black hole and Reissner-Nordström-(anti) de Sitter-like black hole, while the stationary and axially symmetric soltuions describe the Kerr-Newman-like black hole and Kerr-Newman-(anti) de Sitter-like black hole. We utilize the Hamilton-Jacobi formalism to study the shadows of the black holes. Additionally, we investigate the effect of the electric charge and Lorentz-violating parameters on the radius of the shadow reference circle and the distortion parameter. We find that the radius of the reference circle decreases with the Lorentz-violating parameter and the charge parameter, while the distortion parameter increases with the Lorentz-violating parameter and the charge parameter.
@@ I INTRODUCTION
Several quantum gravity theories have been proposed Birrell and Davies (1984); Maldacena (1998); Aharony et al. (2000); Gubser et al. (1998); Alfaro et al. (2000, 2002); Rovelli and Smolin (1990), but it is needed the experiments at the Planck scale ($\sim 10^{19}$ GeV) to observe the quantum effect, this is beyond the ability of current experiments. However, some quantum gravity theories assume that Lorentz symmetry might be broken in the gravitational UV regime which is at our presently accessible low-energy scales. Stemming from the concept of spontaneous Lorentz symmetry breaking (LSB) in string theory Kostelecky and Samuel (1989a); Kostelecky (2004), the Standard Model extension was introduced as a suitable framework for violating Lorentz symmetry Colladay and Kostelecky (1997, 1998); Kostelecky (2004).In the following decades, the Standard Model extension has been widely studied Colladay and Kostelecky (2001); Kostelecky and Lehnert (2001); Kostelecky and Mewes (2001); Colladay and McDonald (2009); Berger et al. (2016); Carroll et al. (1990); Andrianov and Soldati (1995, 1998); Lehnert and Potting (2004); Baeta Scarpelli (2012); Brito et al. (2013).
@@
In 1989, Kostelecky and Samuel introduced the prototype of the bumblebee model Kostelecky and Samuel (1989b, a), a string-inspired framework featuring tensor-induced spontaneous Lorentz symmetry breaking. In the framework of bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B^{\mu}B_{\mu})$ acting on an vector field $B_{\mu}$ Kostelecky (2004); Kostelecky and Tasson (2011). Not just in Minkowski’s spacetime, the bumblebee model can also be explored in Riemann and Riemann-Cartan spacetimes Bailey and Kostelecky (2006); Maluf et al. (2013); Li and Övgün (2020); Bluhm et al. (2008a); Escobar and Martín-Ruiz (2017); Maluf et al. (2015); Hernaski (2014); Carroll et al. (2009).
@@
The pursuit of black hole solutions holds significance for any theory of gravity. In 2017, Casana et al. obtained an exact Schwarzschild-like black hole solution in the bumblebee gravity Casana et al. (2018). Subsequently, researchers have independently proposed exact solutions for traversable wormholes Övgün et al. (2019), Schwarzschild-anti-de Sitter (AdS)-like black holes Maluf and Neves (2021), slowly rotating Kerr-like black holes Ding et al. (2020), and other solutions Santos et al. (2015); Jha and Rahaman (2021); Filho et al. (2023); Xu et al. (2023); Ding et al. (2023) within bumblebee gravity. Moreover, in other scenarios within the framework of the Standard Model Extension, such as when considering the nonminimally coupled Kalb-Ramond field with a nonzero vacuum expectation value, certain black hole solutions have also been obtained Lessa et al. (2020); Kumar et al. (2020); Yang et al. (2023); Duan et al. (2023); Do and Kao (2020); Liu et al. (2024a). And the properties of the blakc holes were also studied Guo et al. (2023); Du et al. (2024); Kuang and Övgün (2022); Oliveira et al. (2019); Liu et al. (2023); Gomes et al. (2020); Kanzi and Sakallı (2019); Güllü and Övgün (2022); Oliveira et al. (2021); Kanzi and Sakallı (2021); Hosseinifar et al. (2024); Liu et al. (2024b, a); Ovgün et al. (2018); Sakallı and Yörük (2023); Mangut et al. (2023); Uniyal et al. (2023).
@@ II EINSTEIN-BUMBLEBEE THEORY
$T^{M}_{\mu\nu}=\frac{1}{\kappa}\left[(1+\gamma b^{2})(2F_{\mu\alpha}F_{\nu}^{\alpha}-\frac{1}{2}g_{\mu\nu}F^{\alpha\beta}F_{\alpha\beta})+\gamma B_{\mu}B_{\nu}F^{\alpha\beta}F_{\alpha\beta}\right].$ (7)
@@ III.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
where $\lambda$ is a constant.This form is same as the potential form of the Higgs field, and is also related to the mass structure of the theory Kostelecky and Samuel (1989b). In this case, the potential $V$ has no contribution to the field equations. Other choices of the potential, such as $V(X)\equiv\frac{\lambda}{2}X^{n}$ ($n\geq 3$), also have no contribution to the field equations. Therefore, the solutions for these choices are consistent with the potential $V(X)=\frac{\lambda}{2}X^{2}$.
@@
$\frac{d}{dr}[r^{2}\phi^{\prime}(r)]$ $=$ $0$ (28)
By simplifying Eqs. (24) and (25), we obtain:
@@
Here $\Sigma$ represents a three-dimensional spacelike region with the induced metric $\gamma_{ij}^{(3)}$, the boundary $\partial\Sigma$ is a two-sphere located at spatial infinity with the induced metric $\gamma_{ij}^{(2)}=r^{2}{\leavevmode\nobreak\ d}\Omega^{2}$, while $n_{\mu}=(1,0,0,0)$ and $\sigma_{\mu}=(0,1,0,0)$ denote the unit normal vectors associated with $\partial\Sigma$ and $\Sigma$, respectively. The solution of $A(r)$ can also be rewritten as
@@
this solution is similar to the Reissner-Nordström (RN) solution. As $\ell$ tends to 0, it recovers to the RN solution. The introduction of Lorentz symmetry breaking distinguishes it from many modified gravity theories. It is evident that the metric functions approach to $A(r)\rightarrow 1$ and $S(r)\rightarrow 1+\ell$ at infinity, indicating that the spacetime is not asymptotically Minkowski.Furthermore, the $g_{tt}$ component of the metric (13) differs from the RN black hole solution only by the modification of the coefficient of the $\frac{Q^{2}}{r^{2}}$ term. However, the electrically charged black hole solutions in many modified gravity theories modify the terms involving both $\frac{M}{r}$ and $\frac{Q^{2}}{r^{2}}$ in the RN metric, or introduce new terms. For example, the charged black hole solutions in Einstein-Gauss-Bonnet gravity Fernandes (2020) and Eddington-inspired Born–Infeld gravity Wei et al. (2015a) modify the terms involving both $\frac{M}{r}$ and $\frac{Q^{2}}{r^{2}}$, while the charged black hole solution in Einstein-Maxwell-aether gravity introduces new term Ding et al. (2015). In the paper Casana et al. (2018), Casana et al. studied the effects of spontaneous Lorentz symmetry breaking and investigated some classic tests, including the advance of Mercury’s perihelion, bending of light, and Shapiro’s time-delay. Furthermore, they have computed several upper bounds, with the most stringent sensitivity of $\ell<6.2\times 10^{-13}$.
The introduction of the nonminimal coupling parameter $\ell$ does have some effects on black hole thermodynamics, such as the black hole temperature, entropy, and so on. However, it does not fundamentally alter the number of possible black hole phases, the phase transition behaviors, and the thermodynamic stability. In Appendix A, we have provided a series of specific analyses and discussions.
Similar to the RN black hole, this solution exhibits two horizons:
@@
From this we can see that, the singularity only appears at $r=0$ (assuming $0<\left|\ell\right|\ll 1$) which is similar to the singularity of the RN black hole.
To further clarify and describe the spacetime structure of the Charged black hole in Bumblebee gravity, we will present the Penrose diagram. By performing a coordinate transformation, we obtain:
${d}{s}^{2}=-(1-R_{+})(1-R_{-}){dt}^{2}+\frac{1}{(1-R_{+})(1-R_{-})}{dR}^{2}+\frac{1}{(1+\ell)}R^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (39)
Figure 1: Penrose diagram of the non-extremal charged black hole in bumblebee gravity.
where $R$ is the parameter after coordinate transformation of $r$ ($r=\sqrt{\frac{1}{1+\ell}}R$) and $R_{\pm}$ are the result of the two horizon radii after the coordinate transformation ($R_{\pm}=\sqrt{\frac{1}{1+\ell}}(M\pm\sqrt{M^{2}-\frac{2(\ell+1)}{\ell+2}Q_{0}^{2}})$). The corresponding Penrose diagrams are illustrated in Fig. 1. It is evident that the Penrose diagram of this solution is identical to that of the RN spacetime. The number and type of horizons of the electrically charged black holes obtained in our paper are indeed consistent with the RN black hole. The speed of light remains unchanged, and the causal structure is unaffected. However, due to Lorentz symmetry breaking in bumblebee gravity, the asymptotic behavior of the resulting spacetime is not Minkowskian. In addition, although the Penrose diagram obtained here is the same as that of the RN black hole, the two-dimensional spheres at the same points on the Penrose diagrams are not the same. Two horizon radii of the electrically charged black hole in bumblebee gravity differ from those of the RN black hole, and there is an additional correction factor 1($\frac{1}{(1+\ell)}$ in front of the $R^{2}d\Omega^{2}$ term compared to the RN black hole. It follows that, although our Penrose diagram is identical to that of the RN black hole after a series of coordinate transformations, the radii of the corresponding two-dimensional spheres in the diagrams are different.
@@ III.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
where $\lambda$ is a Lagrange multiplier Bluhm et al. (2008b). Note that the equation of motion for the Lagrange multiplier ensures the vacuum condition $X=0$, which also ensures that $V=0$ for any field $\lambda$ on-shell. However, unlike the vacuum condition in Case A, $V^{\prime}\neq 0$ is valid when the field $\lambda$ is nonzero. Thus, the potential $V$ contributes to the field equations. The choice $V(X)=\frac{\lambda}{2}X$ effectively freezes motion about the potential minimum and hence permits an efficient extraction of the essential physics Kostelecky and Samuel (1989b). The constraint for the Lagrange multiplier $\lambda$ is providedthe field Eqs (10) and (11) .
@@
An analytical solution is possible if and only if
@@
$A(r)$ $=$ $1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)\Lambda_{e}r^{2},$ (47)
$S(r)$ $=$ $\frac{1+\ell}{1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)\Lambda_{e}r^{2}},$ (48)
@@
Figure 2: Penrose diagram of the non-extremal charged AdS black hole in bumblebee gravity.
As $\ell$ approaches to 0, this solution becomes RN-(A)dS black hole solution. Similarly, we compute the Kretschmann scalar in order to study the singularity. The expression for $K$ is
@@
Through the coordinate transformation mentioned of $r$ ($r=\sqrt{\frac{1}{1+\ell}}R$), we obain
${d}{s}^{2}$ $=$ $-f(R){dt}^{2}+\frac{1}{f(R)}{dR}^{2}+\frac{1}{(1+\ell)}R^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (51)
where $f(R)=(1-\frac{2\sqrt{1+\ell}M}{R}+\frac{2(1+\ell)^{2}Q_{0}^{2}}{(2+\ell)R^{2}}-\frac{1}{3}\Lambda_{e}R^{2})$ then we can get the Penrose diagram of the non-extremal charged AdS black hole, as shown in Fig. 2, which is is identical to that of the classical RN-AdS spacetime. But the radii of the two-dimensional spheres at the same points on the Penrose diagrams are not the same.
@@ IV.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
Ding et al. obtained a slowly rotating black hole solution within bumblebee gravity Ding et al. (2020); Ding and Chen (2021). Following their approach and combining with Newman-Janis algorithm Newman and Janis (1965); Newman et al. (1965); Janis and Newman (1965), we obtained a slowly rotating charged black hole solution within bumblebee gravity. The metric in Boyer-Lindquist coordinates reads:
@@
When $\ell\rightarrow 0$, it recovers to the usual Kerr-Newman metric and when $a\rightarrow 0$, it becomes to metric (13) with Eqs. (30) and (31). By solving $\Delta_{r}=0$, we can obatin the radii of the black hole’s horizon
$r_{\pm}=M\pm\sqrt{M^{2}-\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}}.$ (60)
It becomes evident that the mass and spin parameters of the black hole must satisfy the following condition
$\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}\leq M^{2}.$ (61)
@@
$\Delta_{\mu\nu}=$ $R_{\mu\nu}$ $-\Lambda g_{\mu\nu}-\kappa\left[V^{\prime}\left(2B_{\mu}B_{\nu}+b^{2}g_{\mu\nu}\right)+B_{\mu}^{\ \alpha}B_{\nu\alpha}+Vg_{\mu\nu}\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}\right]$ (62)
@@
$\Delta_{\mu\nu}=\left\{\begin{aligned} \mathcal{O}(a^{2}\ell)&,&\quad\mu\nu=tt,rr,\phi\phi,\theta\theta,r\theta,t\phi\\ 0&,&\quad\text{otherwise}\end{aligned}\right.,$ (63)
At the same time, we derive the equations of motion for the bumblebee field and the electromagnetic field. Only the $r$ component nonvanishing:
@@
Combining the above equations, we find that the rotating solution needs to satisfy the condition $\mathcal{O}(a^{2}\ell)$ being a small quantity. This implies that when considering this solution, $a$ and $\ell$ cannot both take large values simultaneously. Thus, for our solution, we need the condition that $a^{2}\ell$ is small quantities ($\frac{a^{2}\left|\ell\right|}{M^{2}}\ll 1$) for finite $Q_{0}$.
Theoretically, there is no constraint on the spin parameter $a$ and the Lorentz-violating parameter $\ell$. The constraint $a^{2}\left|\ell\right|\ll M^{2}$ or $a^{2}\left|\ell\right|\ll 1$ for $M=1$ on the parameters is imposed to ensure that our analytical solution approximately satisfies the field equations. We will use the analytical solution to study some observables of the black hole shadow, under the condition $a^{2}\left|\ell\right|\ll 1$. This constraint affects the range of the parameter combination $a^{2}\ell$, but it does not affect the individual parameter values. As $\ell$ becomes sufficiently small, $a$ may take larger values and even approach $M$. Conversely, when $a$ is sufficiently small, $\ell$ may take larger values. However, based on previous studies, it has been found that $\ell$ is constrained to a small value Casana et al. (2018), so in astrophysical scenarios, we consider $\ell$ to be sufficiently small, allowing a large range for $a$, which can even approach $M$.
When $\ell$ is sufficiently small, we can consider the extreme black hole case. The extreme black hole condition is approximately given by $\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}=M^{2}$, since our solutions are not exact. For sufficiently small $\ell$, the extreme black hole condition can be approximated as $Q_{0}^{2}+a^{2}=M^{2}.$ From this, we can see that $a^{2}\leq M^{2}$. As long as $\ell$ is sufficiently small, $a^{2}\left|\ell\right|\ll M^{2}$ is still satisfied. In summary, provided that $\ell$ remains sufficiently small, our solution remains valid, even for the near-extremal KN-like black hole case.